OpenStudy (anonymous):

apply the quadratic formula to find the roots of the given function, and then graph the function f(x)=y^2-9 please show work for credit

OpenStudy (anonymous):

OM goodness! :/

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thats what i said

OpenStudy (anonymous):

ha ah.. Hey, are you sure that's written correctly? :/

OpenStudy (anonymous):

usually it's f(x) = x^2 - 9

OpenStudy (anonymous):

not y^2 - 9

OpenStudy (anonymous):

f(x) = x^2 - 9

OpenStudy (anonymous):

ok, soooooo. huh... I guess f(x) = x?? lol

OpenStudy (anonymous):

i am having problems figuring out a b c for the formula

OpenStudy (anonymous):

weird. then it should be f(y) = y^2 -9 because we are dealing with a function of y

OpenStudy (anonymous):

it is changed to x

OpenStudy (anonymous):

f(x) = x^2 - 9

OpenStudy (anonymous):

AND there is no need to use the quadratic because you can factor this! x^2 - 9 = x^2 - 3^2 = (x+3) (x-3) so x is simply plus or minus 3! use the quadratic? for what! lol

OpenStudy (anonymous):

finding the roots

OpenStudy (anonymous):

well, the roots are -3 and 3 but anywyay with the quadratic you want to use (1)x^2 + 0x - 9 a = 1 b = 0 c = -9

OpenStudy (anonymous):

ok now I think I got it from there